A stunt-person is doing a Tarzan-like swing on a rope for a movie. She has a mass of 58.0 kg and the rope is 2.40 m long. She starts with the rope at an angle of 43.0 ? from the vertical and she starts from rest.

A) What is the gravitational potential energy of the stunt-person-and-Earth, compared with the potential energy when she is at the bottom of her swing?

B) How fast will she be moving at the bottom of the swing?

C) How much work does the tension in the ropes do as she swings from the initial position to the bottom?
Answer #1

a)

the gravitational potential energy of the stunt-person-and-Earth, compared with the potential energy when she is at the bottom of her swing is

PE=mgh =58*9.81*2.4*(1-Cos43) =366.85 Joules

b)

By Conservation of energy

(1/2)mV2=366.85

V=sqrt[2*366.85/58]

V=3.5567 m/s

c)

work does the tension in the ropes do as she swings from the initial position to the bottom

W=Fd =0

(Since the rope does not extend any distance)